The domestic airfare for business travel for the current year and the previous year was reported. Below is a sample of 12 flights with their domestic airfares shown for both years. Current Year Previous Year Current Year Previous Year 327 507 324 360 450 525 447 399 525 252 462 417 558 426 624 684 408 498 336 486 600 474 501 543 a. Formulate the hypotheses and test for a significant increase in the mean domestic airfare for business travel for the one-year period. What is the p-value? (to 3 decimals) t-value Degrees of freedom p-value is greater than .20 Using a .05 level of significance, what is your conclusion? We cannot conclude b. What is the sample mean domestic alrfare for business travel for each year? that there has been a significance increase in business travel airfares over the one-year period. (to 2 decimals) Current Year Previous Year c. What is the percentage change in the airfare for the one-year period? (to 2 decimals) 96 (to 2 decimals)
Answer #1

From the given data

a)
Test Statistic, t: 0.4893
Degrees of freedom: 22.0000

P-Value: 0.6853

b)
Current Year: 474.3333
Previous Year: Mean: 453.5

c)

Total Variation: 97143
Explained Variation: 3802.624
Unexplained Variation: 93340.38

R^2 = 3802.624 / 97143 = 0.03914 = 3.91%
Answer: 3.91%